# Experimental validation of a linear momentum and bluff-body model for high-blockage cross-flow turbine arrays

### v1

@@ 2.2 Non-Dimensional Parameters
where $h$ is the dynamic water depth and $w$ is the channel width. As subsequently discussed, because of the array layout, global and local blockage for this array are identical and can be described by a single value. The blockage ratio is varied by holding the array geometry constant and changing the water depth in the flume. However, to experimentally isolate the effects of blockage on array performance while changing the water depth, other non-dimensional flow parameters must be controlled. Here, to increase the blockage ratio, the water depth is decreased. This causes an increase in the Froude number based on channel depth,
@@
where $g$ is the acceleration due to gravity. The array’s proximity to the free surface also changes as depth decreases, represented here by the normalized submergence, $s/h$, where $s$ is the distance between the free surface and the top of the turbine blades. Both $Fr_{h}$ (Consul et al., 2013; Hunt et al., 2020; Ross & Polagye, 2022) and $s/h$ (Birjandi et al., 2013; Kolekar & Banerjee, 2015; Kolekar et al., 2019; Ross & Polagye, 2022) have been shown to influence turbine performance. Therefore, as $h$ decreases, $U_{\infty}$ and $s$ must also decrease to hold $Fr_{h}$ and $s/h$, respectively, constant.
@@
where $\nu$ is the kinematic viscosity. The Reynolds number dependence for turbine performance is well documented (Miller et al., 2018, 2021; Bachant & Wosnik, 2016; Ross & Polagye, 2022; Hunt et al., 2024). Although turbine performance becomes independent of the Reynolds number above a certain threshold, for cross-flow turbines, this is difficult to achieve at laboratory scale without the use of compressed-air wind tunnels (Miller et al., 2018, 2021). To compensate for the decrease in $U_{\infty}$ necessitated by holding $Fr_{h}$ constant, the kinematic viscosity is decreased by changing the water temperature. In this way, $\beta$ can be varied while holding $Fr_{h}$, $s/h$, and $Re_{D}$ constant.
@@
Array performance is measured at blockage ratios ranging from $30.0\%$ to $55.0\%$, which is the widest range of blockages possible given the size of these turbines and flume capabilities. The flume conditions required to achieve each blockage while holding $Fr_{h}$ and $Re_{D}$ constant are summarized in table 2. Across all experiments, the measured $\beta$ are within $1.5\%$ of the target values in table 2, and the measured $Fr_{h}$ and $Re_{D}$ do not deviate more than $5\%$ from the values in table 2. The measured turbulence intensity is ${\sim}2\%$ for all test conditions.
The value of $Re_{D}$ in this study is constrained by the maximum $U_{\infty}$ for the highest blockage ratio. This maximum velocity, in turn, is constrained by the entrainment of air from the free surface into the rotor (i.e., ventilation), which decreases the lift forces from the blades and degrades performance (Birjandi et al., 2013; Young et al., 2017). Consequently, the maximum $U_{\infty}$ for the highest tested blockage ratio is set such that any ventilation occurs beyond the rotation rate corresponding to maximum turbine performance. Since ventilation becomes more likely as $s/h$ decreases, to further limit the risk of ventilation, $s/h$ is maximized at each blockage rather than held constant across all blockages. Array performance at various $s/h$ and $\beta$ is separately evaluated (Appendix A), and confirms that varying $s/h$ has minor effects on performance at these test conditions before the onset of ventilation.
2.3 Array Layout and Control
An overhead view of the array layout is shown in figure 3. The center-to-center spacing between the turbines is ${\sim}1.2D$, and the array is positioned laterally such that the blade-to-blade spacing between adjacent turbines (${\sim}0.2D$; ${\sim}0.9c$) is twice the wall-to-blade spacing (i.e., the walls notionally correspond to symmetry planes in a larger array). The turbines are operated under a counter-rotating, phase-locked scheme, wherein both turbines rotate at the same, constant speed, but in opposite directions (i.e., $\omega_{A}=-\omega_{B}$), with a constant angular phase offset, $\Delta\theta$, between them. The turbines are counter-rotated such that the blades rotate toward the array centerline, which has been shown to augment performance relative to other rotation schemes (Zanforlin & Nishino, 2016; Scherl, 2022; Gauvin-Tremblay & Dumas, 2022). The present experiments are conducted at $\Delta\theta=0^{\circ}$, an operating case in which the lateral forces and torques for a pair of counter-rotating turbines are equal and opposite. A closed-loop controller maintains $\Delta\theta$ to within $1^{\circ}$ of the target value across all experiments.
2.4 Performance Metrics
@@
where $\omega=|\omega_{A}|=|\omega_{B}|$ is the magnitude of the angular velocity. Data are collected at each tip-speed ratio for 60 seconds, and the time series is cropped to an integer number of turbine rotations before performance metrics are calculated.
The efficiency (formally, the coefficient of performance) of the array is equal to the net mechanical power produced by the turbines normalized by the kinetic power in the freestream flow that passes through array’s projected area
@@
where $Q_{A}$ and $Q_{B}$ are the the hydrodynamic torques on Turbines A and B, respectively, and $\rho$ is the density of the working fluid. Structural loads on the array are characterized via the thrust coefficient, given as
@@
where $T_{A}$ and $T_{B}$ are the streamwise forces on Turbines A and B. Cross-flow turbines also experience forces in the cross-stream direction ($L_{A}$ and $L_{B}$ in figure 3), which can be similarly non-dimensionalized as an array lateral force coefficient using an analogous equation to Eq. 7. The measured lateral forces are tangential to this work, but provided as supplementary material.
2.5 Flow Visualization
@@
Figure 4: Schematic of PIV test set including turbines and example flow field data.
@@
Figure 5: Location of PIV field of view within the flume, looking downstream. The black rectangles indicate the projected areas of the turbines. The purple lines indicate the nominal location of the free surface at each blockage ratio.
@@
Two-camera, three-component particle image velocimetry is used to unobtrusively record the near-wake flow field in planes perpendicular to the flow (figure 5). Two cameras (Imager sCMOS 5.5 MPx, maximum 50 fps) with 35 mm and 60 mm lenses, are positioned on either side of the flume at approximately $45^{\circ}$ incidence to the image plane. The cameras are oriented perpendicular to water-filled acrylic prisms shown in figure 5, which are used to reduce refractive distortion from the angled camera positioning relative to the flume. Scheimpflug camera lens adaptors correct any residual focal distortions from angled fields of view (FoV). The FoV of each camera is focused on a domain centered behind one of the turbines and results in an overlapping 42.3$\times$27.7 cm2 data collection region. The resulting FoV (figure 5) is positioned approximately 5.2 cm above the bottom of the flume and approximately 8.75 cm from the side wall to maximize capture of the flow through and around the array while reducing overexposure from laser scattering by solid boundaries, and encompasses the majority of the projected area of Turbine B, the bypass region above Turbine B, and the bypass region between the turbines. Although the FoV does not capture the full array, given the layout symmetry, the flow measured around Turbine B is expected to be descriptive of that near Turbine A. The entire flume is seeded with neutrally buoyant hollow glass spheres with a diameter of $\sim\!10\mu\mathrm{m}$ (Potters Industries Sphericel 110P8), and illumination is provided by a 532 nm Nd:YAG laser (EverGreen 200), which produces a light sheet approximately 2 mm thick.
@@
The near-wake flow field is measured in vertical planes $0.6D$ and $1.5D$ downstream of the turbines’ axes of rotation, with the array moved in the streamwise direction to change the relative location of the vertical plane. We focus our analysis on the plane nearest to the array, at $0.6D$, with the flow fields at $1.5D$ provided in supplementary material. Flow field data in each plane are collected at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ for tip-speed ratios of $\lambda=1.6$, $2.6$, $2.9$, and $3.6$ (all $\beta$) and $\lambda=4.2$ ($\beta=55.0\%$ only). These tip-speed ratios include the “optimal” $\lambda$ corresponding to maximum $C_{P}$ for $\beta=36.7\%$, $45.0\%$, and $55.0\%$, as well as $\lambda$ lower (“underspeed”) and higher (“overspeed”) than the optimal $\lambda$ at these blockage ratios, which are relevant for shedding power above the rated flow speed (Pao & Johnson, 2009). Data acquisition is synchronized with turbine phase ($\theta$), with images collected in $3^{\circ}$ increments at the $\lambda$ corresponding to maximum $C_{P}$ for $\beta=36.7\%$, $45.0\%$, and $55.0\%$, and $12^{\circ}$ increments for all other $\lambda$ cases. A total of 40 frames are collected at each phase and are averaged to reduce measurement noise. All phases are then averaged together to produce representative time-averaged fields.
@@
LaVision Davis 10.2.1 software is used to acquire and process the PIV data. A dual-plane calibration target and self-calibration are employed. The resulting scale factor is 7.97 pix/mm with fitting errors of 0.81 and 0.92 pixels for each camera, respectively. A minimum filter subtracts background noise and an average polynomial filter using an interrogation window of 11$\times$11 pixels minimizes unsteady reflections from the free surface or near-blade regions. Vector calculation employs an initial $128\times 128$ interrogation window with a 50% overlap followed by four sequential reducing passes to $32\times 32$ pixels at a 75% overlap. After vector calculation, three passes of a universal outlier detection median filter (Westerweel & Scarano, 2005) removes vectors with a median normalized residual greater than 1.5 on a filter domain of $7\times 7$ vectors. These vectors were reinstated if their normalized residual was less than 2.5 and alternate correlation peaks had a higher residual. Groups with fewer than five vectors were removed.
3 Experimental Results
3.1 Array Performance
@@
Time-average $C_{P}$ as a function of $\beta$ and $\lambda$ is shown in figure 6. In agreement with prior work in table 1, as $\beta$ increases, the maximum $C_{P}$ increases, the array produces power over a broader range of tip-speed ratios, and the optimal tip-speed ratio increases. Similarly, the time-averaged $C_{T}$ (figure 6) generally increases with $\beta$, particularly at higher blockage ratios. Above $\beta=33.4\%$, maximum $C_{P}$ exceeds the Betz limit, and above $\beta=50.0\%$, maximum $C_{P}$ exceeds unity. Such efficiencies are not violations of energy conservation since the definition of $C_{P}$ (Eq. 6) considers only the kinetic power that passes through the array projected area. This conventional definition neglects the contribution from the fluid’s potential energy, which is appreciably drawn down as $\beta$ and $C_{T}$ increase. As $\beta$ increases, a bounded efficiency metric may resemble the hydraulic efficiency of a hydropower turbine (such as that employed by Takamatsu et al. (1985) or McAdam et al. (2013a)) in which the available power is a function of volumetric flow rate through and the head drop across the rotor. This alternative efficiency metric is explored further in the supplementary material, but is sensitive to the resolution and accuracy of free surface measurements. Here, the conventional definitions of $C_{P}$ and $C_{T}$ facilitate comparison with prior studies, as well as consistency with LMADT—which is formulated using the definition of $C_{T}$ in Eq. 7.
@@
When the array power and thrust coefficients are regressed against $\beta$ at constant tip-speed ratio, approximately linear relationships are revealed, which are shown for a subset of $\lambda$ in figure 7. Below a threshold value of $\lambda$ both $C_{P}$ and $C_{T}$ are independent of the blockage ratio and depend only on the tip-speed ratio, which corresponds to the collapsed regions of the $C_{P}-\lambda$ curves ($\lambda<1.6$) and $C_{T}-\lambda$ curves ($\lambda<1.2$) in figure 6. Similar relations are observed in prior work (Consul et al., 2013; Kolekar & Banerjee, 2015; Badshah et al., 2019). As $\lambda$ increases further, $C_{P}$ and $C_{T}$ vary linearly with the blockage ratio for $\beta\leq 45.0\%$ and moderate $\lambda$, which is emphasized by the linear regressions in figure 7. While similar linear relationships between the blockage ratio, power, and thrust have been previously identified in simulations of oscillating foils (Gauthier et al., 2016), axial-flow turbines (Kinsey & Dumas, 2017; Abutunis & Menta, 2022) and cross-flow turbines (Kinsey & Dumas, 2017), this is the first experimental demonstration of this trend. Finally, as $\beta$ and $\lambda$ increase further, these relationships become non-linear with performance increasingly sensitive to incremental changes in the blockage and tip-speed ratios, which is similar to narrower observations for $C_{P}$ by Gauthier et al. (2016)
3.2 Flow Fields
@@
Figure 8 shows the normalized time-average streamwise velocity field $0.6D$ downstream of the array (i.e., within the near-wake) as a function of $\beta$ and $\lambda$. For each $\beta$ and $\lambda$, the velocity field is characterized by a wake region behind the turbines that is slower than the freestream ($U_{x}/U_{\infty}<1$, indicated by blue hues) and a bypass region above, below, and between the turbines where the flow is accelerated ($U_{x}/U_{\infty}>1$, indicated by red hues). We note that the bypass flow region between Turbine B and the flume side wall lies outside of the field of view, and we attribute the persistent region of decelerated flow below the turbine projected area at all $\lambda$ to obstruction of the flow by the test rig (figure 3).
@@
Figure 9: Time-averaged normalized streamwise velocity profiles measured $0.6D$ downstream of the array for various $\beta$ at 9 $\lambda=1.6$, 9 $\lambda=2.6$, and 9 $\lambda=3.6$. The profiles correspond to the average velocity along the middle $70\%$ of the turbine blade span. Vertical gray rectangles correspond to the bypass region and delineate the relative turbine position and wake flow region (white).
Trends in the near-wake velocity with $\beta$ and $\lambda$ are more apparent when visualized as two-dimensional velocity profiles of the average velocity field over the middle $70\%$ of the blade span at each $y/D$ position, as shown in figure 9 for $\lambda=1.6$, $2.6$, and $3.6$ as a function of $\beta$. Ripples in the velocity profiles in the wake region are attributed to the passage of vortices shed from the blades during their downstream sweep. At $\lambda=1.6$ (figure 9), where array thrust varies weakly with $\beta$ (figure 7), $U_{x}/U_{\infty}$ increases slightly with $\beta$ throughout the entire profile, although the general shape of the velocity profile is constant. As $\lambda$ increases (figures 9 and 9), the turbines impart greater resistance on the flow and more fluid is diverted around them, which causes $U_{x}/U_{\infty}$ to increase in the bypass region (i.e., the gray shaded areas in figure 9) at all $\beta$. Similarly, as the blockage ratio increases at constant $\lambda$, the bypass velocity increases. $U_{x}/U_{\infty}$ at a given $\lambda$ also tends to increase in the wake region with $\beta$ since confinement increases the flow through the turbines. However, the increase in wake velocity is less pronounced than the corresponding increase in bypass velocity since momentum is extracted from the flow that passes through the turbines. Although the maximum bypass velocity (which occurs at the array centerline) increases with both the blockage ratio and the tip-speed ratio, $U_{x}/U_{\infty}$ decreases more rapidly away from the centerline at higher $\lambda$, as indicated by the narrowing of the peaks in the velocity profiles in figure 9 at all $\beta$ as $\lambda$ increases. Conversely, an increase in $\beta$ at constant $\lambda$ results in a more gradual reduction of $U_{x}/U_{\infty}$ away from the array centerline, which implies increased turbulent mixing between the bypass flow and the core flow through the turbine. As discussed by Nishino & Willden (2012a), this mixing between the bypass and core flows provides a secondary enhancement of power with increasing blockage by increasing the pressure drop across the rotor as a consequence of increasing velocity in the near-wake.
3.3 Free Surface Deformation

### v2

@@ II.2 Non-Dimensional Parameters
where $h$ is the dynamic water depth and $w$ is the channel width. As subsequently discussed, because of the array layout, global and local blockage for this array are identical and can be described by a single value. The blockage ratio is varied by holding the array geometry constant and changing the water depth in the flume. To isolate blockage effects, other non-dimensional flow parameters must be controlled as the depth changes. For example, decreasing water depth (increasing $\beta$) increases the depth-based Froude number,
@@
where $g$ is the acceleration due to gravity. The array’s proximity to the free surface also changes as depth decreases, represented here by the normalized submergence, $s/h$, where $s$ is the distance between the free surface and the top of the turbine blades. Both $Fr_{h}$ [35, 49, 45] and $s/h$ [39, 38, 50, 45] have been shown to influence turbine performance. Therefore, as $h$ decreases, $U_{\infty}$ and $s$ must also decrease to hold $Fr_{h}$ and $s/h$, respectively, constant.
@@
where $\nu$ is the kinematic viscosity. The dependence of turbine performance on the Reynolds number is well documented [51, 52, 53, 45, 54]. Although turbine performance becomes independent of the Reynolds number above a certain threshold, for cross-flow turbines, this is difficult to achieve at laboratory scale without a compressed-air wind tunnel [51, 52] or relatively large tow tank [53]. To compensate for the decrease in $U_{\infty}$, the kinematic viscosity is decreased by changing the water temperature. In this way, $\beta$ can be varied while holding $Fr_{h}$, $s/h$, and $Re_{D}$ constant.
@@
Array performance is measured at blockage ratios from $30.0\%$ to $55.0\%$, which is the widest range possible given the size of these turbines and flume capabilities. The inflow conditions required to achieve each blockage while holding $Fr_{h}$ and $Re_{D}$ constant are summarized in Table 2. Across all experiments, the measured $\beta$ are within $1.5\%$ of the target values in Table 2, and the measured $Fr_{h}$ and $Re_{D}$ do not deviate more than $5\%$ from the target values.
The value of $Re_{D}$ in this study is constrained by the maximum $U_{\infty}$ for the highest blockage ratio. This maximum velocity, in turn, is constrained by the entrainment of air from the free surface into the rotor (i.e., ventilation), which decreases lift on the blades and degrades performance [39, 55]. Consequently, the maximum $U_{\infty}$ for the highest tested blockage ratio is set such that any ventilation occurs beyond the rotation rate corresponding to maximum turbine performance. To further limit the risk of ventilation, $s/h$ is maximized at each blockage rather than held constant across all blockages, since ventilation becomes more likely as $s/h$ decreases. Array performance at various $s/h$ and $\beta$ is separately evaluated (Appendix A) and confirms that varying $s/h$ has minor effects on performance at these test conditions before the onset of ventilation.
II.3 Array Layout and Control
An overhead view of the array layout is shown in Figure 3. The turbines are positioned laterally such that the center-to-center spacing is ${\sim}1.2D$ and the blade-to-blade spacing between adjacent turbines (${\sim}0.2D$; ${\sim}0.9c$) is twice the wall-to-blade spacing (i.e., the walls notionally correspond to symmetry planes in a larger array). The turbines are operated under a counter-rotating, phase-locked scheme, wherein both turbines rotate at the same, constant speed, but in opposite directions (i.e., $\omega_{A}=-\omega_{B}$), with zero phase offset between them. The turbines are counter-rotated such that the blades rotate toward the channel centerline, which has been shown to augment performance relative to other rotation schemes [56, 57, 58].
II.4 Performance Metrics
@@
where $\omega=|\omega_{A}|=|\omega_{B}|$ is the magnitude of the angular velocity. Data are collected at each tip-speed ratio for 60 seconds, and the time series is cropped to an integer number of turbine rotations.
The array efficiency (power coefficient) is the ratio of net mechanical power produced by the turbines to the kinetic power in the freestream flow that passes through array’s projected area
@@
where $Q_{A}$ and $Q_{B}$ are the the hydrodynamic torques on the turbines and $\rho$ is the density of the working fluid. Structural loads on the array are characterized via the thrust coefficient
@@
where $T_{A}$ and $T_{B}$ are the streamwise forces on the turbines. Cross-flow turbines also experience forces in the cross-stream direction ($L_{A}$ and $L_{B}$ in Figure 3), which can be similarly non-dimensionalized as an array lateral force coefficient using an analogous equation to Eq. 7. The lateral forces are tangential to this work, but provided as supplementary material [59].
II.5 Flow Visualization
@@
Figure 4: Schematic of the PIV test setup including turbines and example flow field data.
@@
Figure 5: Location of PIV field of view within the flume, looking downstream. The lateral and vertical coordinates for the field of view are defined with respect to the center of Turbine B. The black rectangles indicate the projected areas of the turbines. The purple lines indicate the nominal location of the free surface at each blockage ratio.
@@
Two-camera, three-component particle image velocimetry is used to measure the near-wake flow field in planes perpendicular to the flow (Figure 5). Two cameras (Imager sCMOS 5.5 MPx, maximum 50 fps) with 35 mm and 60 mm lenses, are positioned on either side of the flume at approximately $45^{\circ}$ incidence to the image plane. The cameras are oriented perpendicular to water-filled acrylic prisms shown in Figure 5, which are used to reduce refractive distortion from the angled camera positioning relative to the flume. Scheimpflug camera lens adaptors correct any residual focal distortions from angled fields of view (FoV). The FoV of each camera is focused on a domain centered behind one of the turbines and results in an overlapping 42.3$\times$27.7 cm2 data collection region. The resulting FoV (Figure 5) is positioned approximately 5.2 cm above the bottom of the flume and approximately 8.75 cm from the side wall to maximize capture of the flow through and around the array while reducing overexposure from laser scattering by solid boundaries. The FoV encompasses the majority of the projected area of Turbine B, the bypass region above Turbine B, and the bypass region between the turbines. Although the FoV does not capture the full array, given the layout symmetry, the flow measured around Turbine B is expected to be descriptive of that near Turbine A. The entire flume is seeded with neutrally buoyant hollow glass spheres with a diameter of $\sim\!10\mu\mathrm{m}$ (Potters Industries Sphericel 110P8), and illumination is provided by a 532 nm Nd:YAG laser (EverGreen 200), which produces a light sheet approximately 2 mm thick.
@@
The near-wake flow field is measured in vertical planes $0.6D$ and $1.5D$ downstream of the turbines’ axes of rotation, with the array moved in the streamwise direction to change the relative location of the vertical plane. We focus our analysis on the plane nearest to the array at $0.6D$, with the flow fields at $1.5D$ provided in supplementary material [59]. Flow field data in each plane are collected at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ for tip-speed ratios of $\lambda=1.6$, $2.6$, $2.9$, and $3.6$ (all $\beta$) and $\lambda=4.2$ ($\beta=55.0\%$ only). These tip-speed ratios include the “optimal” $\lambda$ corresponding to maximum $C_{P}$ for $\beta=36.7\%$, $45.0\%$, and $55.0\%$, as well as $\lambda$ lower (“underspeed”) and higher (“overspeed”) than the optimal $\lambda$ at these blockage ratios, which are relevant in field deployments for regulating turbine power above the rated flow speed [60]. Data acquisition is synchronized with turbine phase ($\theta$), with images collected in $3^{\circ}$ increments at the $\lambda$ corresponding to maximum $C_{P}$ for $\beta=36.7\%$, $45.0\%$, and $55.0\%$, and $12^{\circ}$ increments for all other $\lambda$ cases. A total of 40 frames are collected at each phase and are averaged to reduce measurement noise. All phases are then averaged together to produce representative time-averaged fields.
@@
LaVision Davis 10.2.1 software is used to acquire and process the PIV data. A dual-plane calibration target and self-calibration are employed. The resulting scale factor is 7.97 pix/mm with fitting errors of 0.81 and 0.92 pixels for each camera, respectively. A minimum filter subtracts background noise and an average polynomial filter using an interrogation window of 11$\times$11 pixels minimizes unsteady reflections from the free surface or near-blade regions. Vector calculation employs an initial $128\times 128$ interrogation window with a 50% overlap followed by four sequential reducing passes to $32\times 32$ pixels at a 75% overlap. After vector calculation, three passes of a universal outlier detection median filter [61] removes vectors with a median normalized residual greater than 1.5 on a filter domain of $7\times 7$ vectors. These vectors are reinstated if their normalized residual is less than 2.5 and alternate correlation peaks have a higher residual. Groups with fewer than five vectors are removed.
III Experimental Results
III.1 Array Performance
@@
Time-average $C_{P}$ as a function of $\beta$ and $\lambda$ is shown in Figure 6. As for prior studies (Table 1), as $\beta$ increases, the maximum $C_{P}$ increases, the array produces power over a broader range of tip-speed ratios, and the optimal tip-speed ratio increases. Similarly, the time-averaged $C_{T}$ (Figure 6) generally increases with $\beta$, particularly at higher blockage ratios. Above $\beta=33.4\%$, maximum $C_{P}$ exceeds the Betz limit, and above $\beta=50.0\%$, maximum $C_{P}$ exceeds unity. Such efficiencies are not violations of energy conservation since the definition of $C_{P}$ in Eq. 6 considers only the kinetic power passing through the array projected area. This neglects the fluid’s potential energy, which is appreciably drawn down as $\beta$ and $C_{T}$ increase. As $\beta$ increases, a unity-bounded efficiency metric may resemble the hydraulic efficiency of a hydropower turbine (e.g., metrics employed by Takamatsu et al. [34] or McAdam et al. [36]) in which the available power is a function of volumetric flow rate through the channel and the head drop across the array. This alternative efficiency metric is explored further in the supplementary material [59], but is sensitive to the resolution and accuracy of free surface measurements. Here, the conventional definitions of $C_{P}$ and $C_{T}$ facilitate comparison with prior studies, as well as consistency with LMADT—which is formulated using the definition of $C_{T}$ in Eq. 7.
@@
When the array power and thrust coefficients are regressed against $\beta$ at constant tip-speed ratio, approximately linear relationships are revealed, which are shown for a subset of $\lambda$ in Figure 7. Below a threshold value of $\lambda$, both $C_{P}$ and $C_{T}$ are independent of the blockage ratio and depend only on the tip-speed ratio, which corresponds to the collapsed regions of the $C_{P}-\lambda$ curves ($\lambda<1.6$) and $C_{T}-\lambda$ curves ($\lambda<1.2$) in Figure 6. Similar relations are observed in prior work for both cross-flow turbines[35] and axial-flow turbines [38, 42]. As $\lambda$ increases further, $C_{P}$ and $C_{T}$ vary linearly with the blockage ratio for $\beta\leq 45.0\%$ and moderate $\lambda$, which is emphasized by the linear regressions in Figure 7. While similar linear relationships between the blockage ratio, power, and thrust have been previously identified in simulations of oscillating foils [16], axial-flow turbines [17, 19] and cross-flow turbines [17], this is the first experimental demonstration of this trend. Finally, as $\beta$ and $\lambda$ increase further, these relationships become non-linear with performance increasingly sensitive to incremental changes in the blockage and tip-speed ratios, which is similar to narrower observations for $C_{P}$ by Gauthier et al. [16].
III.2 Flow Fields
@@
Figure 8 shows the normalized time-average streamwise velocity field $0.6D$ downstream of the array (i.e., within the near-wake) as a function of $\beta$ and $\lambda$. The velocity fields are characterized by a wake region behind the turbines ($U_{x}/U_{\infty}<1$, indicated by blue hues) and a bypass region with accelerated flow above, below, and between the turbines ($U_{x}/U_{\infty}>1$, indicated by red hues). The bypass flow region between Turbine B and the flume wall lies outside of the field of view, and we attribute the persistent region of decelerated flow below Turbine B at all $\lambda$ to the test rig wake (Figure 3).
@@
Figure 9: Time-averaged normalized streamwise velocity profiles measured $0.6D$ downstream of the array for various $\beta$ at 9 $\lambda=1.6$, 9 $\lambda=2.6$, and 9 $\lambda=3.6$. The profiles correspond to the average velocity along the middle $70\%$ of the turbine blade span. Vertical gray rectangles correspond to the bypass region.
Trends in the evolution of the near-wake with $\beta$ and $\lambda$ are apparent in two-dimensional velocity profiles of the average streamwise velocity at each $y/D$ (Figure 9). These profiles are calculated over the middle $70\%$ of the blade span to exclude edge effects from the blade-end struts. Ripples in the wake region of these profiles are attributed to the passage of vortices shed from the blades during their downstream sweep. At $\lambda=1.6$ (Figure 9), where array thrust varies weakly with $\beta$ (Figure 7), $U_{x}/U_{\infty}$ increases slightly with $\beta$ across the lateral profile. As $\lambda$ increases (Figures 9 and 9), the turbines impart greater resistance on the flow and more fluid is diverted around them, causing $U_{x}/U_{\infty}$ to increase in the bypass region for all $\beta$. Similarly, at a given $\lambda$, the bypass velocity increases with $\beta$. $U_{x}/U_{\infty}$ in the wake region also tends to increase with $\beta$ at a given $\lambda$ since confinement increases the flow through the turbines. However, increases in the wake velocity with $\beta$ are less pronounced than the corresponding increases in bypass velocity due to momentum extraction by the array from the core flow through the turbines. Although the maximum bypass velocity at the array centerline increases with both the $\beta$ and $\lambda$, $U_{x}/U_{\infty}$ decreases more rapidly away from this midpoint at higher $\lambda$. This is indicated by the narrower peaks in the velocity profiles as $\lambda$ increases at a given $\beta$. Conversely, an increase in $\beta$ at constant $\lambda$ results in a more gradual reduction of $U_{x}/U_{\infty}$ away from the array centerline, which implies increased turbulent mixing between the bypass flow and flow through the turbine. As discussed by Nishino and Willden [10], mixing between the bypass and core flows provides a secondary power enhancement with increasing blockage by increasing the pressure drop across the rotor as a consequence of elevated near-wake velocity.
III.3 Free Surface Deformation